OpenStudy (yash2651995):

integrate [ sqrt (tanx) + sqrt (cotx) ] dx limits of integration: lower limit = 0 upper limit = pi/4 prove: it'll be equal to sqrt(2) . (pi /2) need hints. cant get any idea :(

OpenStudy (anonymous):

As i answered yesterday one of the questions, you need to think of conversion here, in other words thinking of this trig function in terms of 1/(1+x^4)

OpenStudy (anonymous):

you need to use u substitution

OpenStudy (anonymous):

this is the answer for integral of sqrt (1/4)*sqrt(2)*(2*sqrt(sin(x)/cos(x))*cos(x)*arccos(cos(x)-sin(x))+sqrt(cos(x)*sin(x))*ln(2)-2*sqrt(cos(x)*sin(x))*ln(cos(x)*sqrt(2)+2*sqrt(sin(x)/cos(x))*cos(x)+sqrt(2)*sin(x)))/sqrt(cos(x)*sin(x))

OpenStudy (yash2651995):

i tried using properties of definate integrals... as i saw pi/4 in limits... and tan (pi/4) = 1 = cot (pi/4) but failed.. tried to convert it in sin and cos stuck again..

OpenStudy (anonymous):

\[\int\limits \sqrt{\tan(x)}dx\]

OpenStudy (anonymous):

i would recommend that you expand the taylor series of each function

OpenStudy (anonymous):

no Taylor series for sqrt tan(x), we need to search for different expansion method of any series...

OpenStudy (anonymous):

is your teacher mad with his wife or something, to fulfill his anger with such an integral...!!????

OpenStudy (anonymous):

i will have to use the whole BW of openstudy to type the solution for this problem...

OpenStudy (yash2651995):

there would be some easy solution...

OpenStudy (anonymous):

let u = tanx, du = sec^xdx, do the same thing for cot(x)

OpenStudy (yash2651995):

without expantion because we havent bee taught any expantion of tan or any such trignometric functions at school.. i guess properties of integral must simplyfy things.. i get sqrt (2) . (sin x + cos x) / sqrt (sin2x)

OpenStudy (yash2651995):

now using properties.. sqrt 2 (sin y+ cos y)/ sqrt (cos 2 x) y = pi/4 - x

OpenStudy (yash2651995):

now at this step i am stuck.. i dont know how to convert sin x + cos x as a derivative of denominator or sumthing like that..

OpenStudy (anonymous):

well even if you use tanx^2=1-csc^2 property, you will gave to take the 4rth root...!!!!

OpenStudy (anonymous):

actually let me try a different method...

OpenStudy (anonymous):

tanx = sinx/cosx

OpenStudy (yash2651995):

ohkay . . .

OpenStudy (anonymous):

cotx=cosx/sinx

OpenStudy (anonymous):

but we will take the root of them

OpenStudy (anonymous):

did you try to simplify this expression

OpenStudy (mimi_x3):

\[\int\limits \sqrt{tanx}+\sqrt{cotx}dx =>\int\limits\frac{sinx+cosx}{\sqrt{sinxcosx}} => \] \[u=sinx-cosx => du=(sinx+cosx)dx\] \[1-2\sin2x=u^2=> \sin2x=(1-u^2)\] \[\int\limits\limits\frac{\sin+cosx}{\sqrt{1/2\sin2x}} =>\sqrt{2} \int\limits\limits\frac{sinx+cosx}{\sqrt{\sin2x}} =>\sqrt{2}\int\limits\limits\frac{1}{\sqrt{1-u^{2}}} du\]

OpenStudy (yash2651995):

^WOW O-O

OpenStudy (yash2651995):

mathsmind, i wanted to do the first few steps... rest of steps were mystery to me... thanks! mimi_x3

OpenStudy (anonymous):

all the best

OpenStudy (inkyvoyd):

MIWIWIWIWIWIWIWIWIWIWI